Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form?
Answer by Maths68(1474)

Points (x,y)(-1,5)
Slope = m = 2
Equation of the straight line point-slope form
m=(y2-y1)/(x2-x1)
Put the values of m, x and y in above equation
2=(y-5)/(x-(-1))
2=(y-5)/(x+1)
2(x+1))=y-5
2x+2=y-5
2x=y-5-2
2x=y-7
y-7=2x
y=2x+7
Above equation is the equation of the line in standard form (y=mx+b)